# Geodesics in the Deep Linear Network

### v1

@@ 3.4.1. Riemannian Geometry Background
Let $(\mathcal{M},g)$ and ($\mathcal{N},h)$ be linked by a differentiable map $\pi:\mathcal{M}\to\mathcal{N}$. Then, there exists a direct sum decomposition of the tangent space at any point $m\in\mathcal{M}$ via the differential of $\phi$.
(3.51) $T_{m}\mathcal{M}=\ker D\pi[m]\oplus(\ker D\pi[m])^{\perp}=\mathcal{V}_{m}\oplus\mathcal{H}_{m},$
@@
Let $(\mathcal{M},g)$ and $(\mathcal{N},h)$ be two Riemannian manifolds and let $T_{m}\mathcal{M}=\mathcal{V}_{m}\oplus\mathcal{H}_{m}$ be the direct sum decomposition of the tangent space at $m\in\mathcal{M}$. A mapping $\pi:\mathcal{M}\to\mathcal{N}$ is a Riemannian submersion if and only if the differential $D\pi[m]:T_{m}\mathcal{M}\to T_{\pi(m)}\mathcal{N}$ satisfies the following for all $m\in\mathcal{M}$:
@@
A core fact underlying [5, Section 4] and Theorem 2.3 is that understanding the horizontal space of a submersion allows us to identify geodesics that are preserved under the submersion.
@@
Let $(\mathcal{M},g)$ and $(\mathcal{N},h)$ be Riemannian manifolds and $\phi:\mathcal{M}\to\mathcal{N}$ be a Riemannian submersion. Let $x:[0,1]\to\mathcal{M}$ be a geodesic between $A,B\in\mathcal{M}$ in the geometry $(\mathcal{M},g)$ so that $x^{\prime}(0)\in\mathcal{H}_{A}$ i.e. is horizontal. Then,
@@
    $\phi\circ x$ is a geodesic in $(\mathcal{N},h)$ between $\phi(A)$ and $\phi(B)$ and is the same length as $x$.
Thus, if we are able to classify horizontal geodesics in $(\mathcal{M},\iota)$, Theorem 3.9 will automatically give explicit formulas for geodesics in $(\mathfrak{M}_{d},g^{N})$ through $\phi$, the submersion. The simplest class of geodesics in $(\mathcal{M},\iota)$ are straight lines. Thus, we begin with a proof of Lemma 2.2 that characterizes the straight lines in $\mathcal{M}$.
@@
We wish to find the condition that $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$ is in $\mathcal{M}$ for all $t\in(0,1)$ if $\mathbf{A},\mathbf{B}\in\mathcal{M}$. Fix $p\in\{1,2,\ldots,N-1\}$. We would like to show that
@@
where $W_{p}(t)=(1-t)W_{p}(0)+tW_{p}(1)$ for all $p$. Expanding the algebra out and using that $W_{p}(0)=A_{p}$ and $W_{p}(1)=B_{p}$ are both in the balanced manifold, we find that $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in(0,1)$ if and only if
@@
The sufficient condition ($R_{p+1}=R_{p-1}$ for all $1\leq p\leq N-1$) clearly results in this condition being satisfied for all $t\in(0,1)$. From here, it is a simple rewriting in terms of $M$ to get the final equivalent condition:
(3.56) $R_{p+1}-R_{p-1}$ $=-M(Q_{p+1}-Q_{p-1})^{T}M^{-T}.$
(3.57) $(R_{p+1}-R_{p-1})M^{T}$ $=-M(Q_{p+1}-Q_{p-1})^{T}.$
The structure demonstrates that the difference
(3.58) $R_{p+1}-R_{p-1}$
behaves like a skew symmetric matrix after changing into coordinates given by $M$ (or essentially $R_{p}$). ∎
We now state an orthonormal basis for the horizontal space in $\mathcal{M}$ computed in [10, Section 6] and [11, Theorem 34]. Let
@@
so that $\mathbf{M}=\xi(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{1})$ and $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{d})$. Furthermore, let $q_{p,i}$ denote the $i$th column of $Q_{p}$. Then, define the following special rank one elements of $\mathbb{M}_{d}^{N}$:
(3.60) $\begin{cases}\mathbf{l}^{k}=\left(l^{k}_{N},l^{k}_{N-1},\ldots,l^{k}_{1}\right),\\ l^{k}_{s}=\frac{1}{\sqrt{N}}q_{s,k}q_{s-1,k}^{T},\end{cases}$
(3.61) $\begin{cases}\mathbf{u}^{k,l,0}=\left(u_{N}^{k,l,0},u_{N-1}^{k,l,0},\ldots,u_{1}^{k,l,0}\right),\\ u^{k,l,0}_{s}=\sqrt{\frac{\lambda_{k}^{2}-\lambda_{l}^{2}}{\lambda_{k}^{2N}-\lambda_{l}^{2N}}}\lambda_{k}^{s-1}\lambda_{l}^{N-s}q_{s,l}q_{s-1,k}^{T},\end{cases}$
and
(3.62) $\begin{cases}\mathbf{u}^{k,l,N}=\left(u_{N}^{k,l,N},u_{N-1}^{k,l,N},\ldots,u_{1}^{k,l,N}\right),\\ u^{k,l,N}_{s}=\sqrt{\frac{\lambda_{k}^{2}-\lambda_{l}^{2}}{\lambda_{k}^{2N}-\lambda_{l}^{2N}}}\lambda_{k}^{N-s}\lambda_{l}^{s-1}q_{s,k}q_{s-1,l}^{T}.\end{cases}$
Lemma 3.10 (Orthonormal Basis for Horizontal Space in $\mathcal{M}$).
(3.63) $\begin{cases}\mathbf{l}=\{l_{k}|1\leq k\leq d\},\\ \mathbf{u}^{0}=\{u^{k,l,0}|1\leq k<l\leq d\},\\ \mathbf{u}^{N}=\{u^{k,l,N}|1\leq k<l\leq d\}\end{cases}$
form a basis for $\mathcal{H}_{\mathbf{M}}=(\ker\phi^{*}[\mathbf{M}])^{\perp}$.
Remark 3.11.
Note a matrix $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{M}}$ (span of $\mathbf{l}$, $\mathbf{u}^{k,l,0}$, and $\mathbf{u}^{k,l,N}$) if and only if there exists a set of coefficients $\alpha\in\mathbb{M}_{d}$ such that
@@
for all $1\leq p\leq N$.
Now that we understand both the straight lines and horizontal space in $\mathcal{M}$, we can use $\phi$ and Theorem 3.9 to construct new explicit formulas for geodesics on $\mathfrak{M}_{d}$.
@@
where $\xi$ is the map given in Equation 1.5 since $\mathbf{A}\in\mathcal{M}$. Clearly, $\phi(\mathbf{A})=A$. Furthermore, Lemma 3.10 gives a set of orthonormal basis vectors for the horizontal space $\mathcal{H}_{\mathbf{A}}$:
(3.68) $l^{k}_{s}$ $=\frac{1}{\sqrt{N}}\begin{cases}u_{k}e_{k}^{T}&s=N\\ e_{k}v_{k}^{T}&s=1\\ e_{k}e_{k}^{T}&\text{o/w}\end{cases},$
(3.69) $u^{k,l,0}_{s}$ $=\sqrt{\frac{\sigma_{k}^{\frac{2}{N}}-\sigma_{l}^{\frac{2}{N}}}{\sigma_{k}^{2}-\sigma_{l}^{2}}}\sigma^{\frac{s-1}{N}}\sigma_{l}^{\frac{N-s}{N}}\begin{cases}u_{l}e_{k}^{T}&s=N\\ e_{l}v_{k}^{T}&s=1\\ e_{l}e_{k}^{T}&\text{o/w}\end{cases},$
(3.70) $u^{k,l,0}_{s}$ $=\sqrt{\frac{\sigma_{k}^{\frac{2}{N}}-\sigma_{l}^{\frac{2}{N}}}{\sigma_{k}^{2}-\sigma_{l}^{2}}}\sigma^{\frac{N-s}{N}}\sigma_{l}^{\frac{s-1}{N}}\begin{cases}u_{k}e_{l}^{T}&s=N\\ e_{k}v_{l}^{T}&s=1\\ e_{k}e_{l}^{T}&\text{o/w}\end{cases}.$
By Remark 3.11, $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{A}}$ if and only if there exists $\alpha\in\mathbb{M}_{d}$ such that
@@
It is immediately obvious that $\mathbf{A}$ is in its own horizontal space $\mathcal{H}_{\mathbf{A}}$ using the coefficient matrix $\alpha=\Sigma^{\frac{1}{N}}$.
@@
where $Q=U^{T}\tilde{U}=V^{T}\tilde{V}$ by assumption. Clearly, $\phi(\mathbf{B})=B$. Furthermore, $\mathbf{A}$ and $\mathbf{B}$ satisfy the sufficient condition of Lemma 2.2 since $Q^{T}=\tilde{U}^{T}U=\tilde{V}^{T}V$ and $R_{p}$ is constant for all $p$. Thus, the line $(1-t)\mathbf{A}+t\mathbf{B}$ is indeed the geodesic between $\mathbf{A}$ and $\mathbf{B}$ in $(\mathcal{M},\iota)$.
Now we will show $x^{\prime}(0)$ is horizontal for the chosen $\mathbf{A}$ and $\mathbf{B}$. We have already shown that $\mathbf{A}$ is in $\mathcal{H}_{\mathbf{A}}$. By Equation 3.66, it is now sufficient to show that $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$ as well. We thus return to solving the system of equations in Remark 3.11 with $\mathbf{Z}=\mathbf{B}$:
@@
(3.77) $\alpha V^{T}$ $=Q\tilde{\Sigma}^{\frac{1}{N}}\tilde{V}^{T},$
and has the obvious solution given by the middle equations of $\alpha=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T}$ which is consistent with the others by definition of $Q$. Thus, $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$.
Since we have shown that both $\mathbf{B}$ and $\mathbf{A}$ are in $\mathcal{H}_{\mathbf{A}}$ and that the geodesic linking them is a straight line, it follows from Equation 3.66 that $x^{\prime}(0)$ is horizontal. Applying Theorem 3.9, we conclude that $\phi\circ x$ is a geodesic in $(\mathfrak{M}_{d},g^{N})$ of the same length as $x$. The explicit formula comes directly from applying $\phi$ to $x$ between our parameterized $\mathbf{A}$ and $\mathbf{B}$. ∎
@@ 4. Open Problems
The main follow up work is (1) solving the general case geodesic equations for arbitrary endpoints and (2) generalizing geodesic results to low rank (rank <d<d) matrices, not just $\mathcal{M}$, as the Riemannian metric and submersion structure exists in all ranks [3, 10, 11]. Once geodesics and distances are understood in the geometries underlying the DLN, one can consider reconnecting them to the vector field given by the gradient flow to make conclusions about training dynamics, similar to [8, Section 3.3].
@@ Acknowledgements
Appendix A Informal Note on Equation 2.11
The geodesic in Equation 2.11 also appears in another geometry [4] when the endpoints are positive definite diagonal matrices. In particular, we endow $\mathbb{P}_{d}$ with the trace metric: at a matrix $P\in\mathbb{P}_{d}$ and tangent vector $Z\in T_{P}\mathbb{P}_{d}$, we let
(A.1) $\|Z\|_{\operatorname{Tr}}^{2}=\operatorname{Tr}(P^{-1}ZP^{-1}Z).$
Consider $P$ positive definite and diagonal, and let $Z$ commute with $P$ (e.g. also diagonal). By the properties of trace,
(A.2) $\|Z\|_{\operatorname{Tr}}^{2}=\operatorname{Tr}(ZP^{-2}Z).$
Between any matrices $P_{0},P_{1}\in\mathbb{P}_{d}$, the geodesic is exactly $P_{0}^{1-t}P_{1}^{t}$. Interestingly, $g^{N}$ as $N\to\infty$ coincides with the trace metric under these assumptions. Consider a variant of $\mathcal{A}_{N,\cdot}$ with an additional $\frac{1}{N}$ factor in front so that a $N\to\infty$ limit exists. Then, indeed, if $P$ is diagonal and commutes with $Z$,
(A.3) $\mathcal{A}_{N,P}(Z)=\frac{1}{N}\sum_{p=1}^{N}(P^{2})^{\frac{N-p}{N}}Z(P^{2})^{\frac{p-1}{N}}=(P^{2})^{\frac{N-1}{N}}Z.$
The limit as $N\to\infty$ is just $P^{2}Z$. Thus, $\mathcal{A}^{-1}_{\infty,P}=P^{-2}Z$ and we conclude that
(A.4) $\|Z\|_{g^{\infty}}=\operatorname{Tr}(Z\mathcal{A}^{-1}_{\infty,P}Z)=\operatorname{Tr}(ZP^{-2}Z)=\|Z\|_{\operatorname{Tr}}.$
More generally, if $X\in\mathfrak{M}_{d}$ is symmetric, for all $Z\in T_{X}\mathfrak{M}_{d}$ where $X$ and $Z$ commute, we have that this agreement between $g^{\infty}$ and the trace metric. Since along the geodesic $X(t)$ between $A$ and $B$ $\dot{X}\in T_{X}\mathfrak{M}_{d}$ is always diagonal, we have agreement between the trace metric and DLN metric for the entire geodesic. Contrast this with the smallest nontrivial depth, $N=2$, where if $X$ is symmetric we know already that the DLN metric coincides with the Bures Wasserstein metric for symmetric tangent vectors.

### v2

@@ 3.4.1. Riemannian Geometry Background
Let $(\mathcal{X},g)$ and ($\mathcal{Y},h)$ be linked by a differentiable map $\pi:\mathcal{X}\to\mathcal{Y}$. Then, there exists a direct sum decomposition of the tangent space at any point $m\in\mathcal{X}$ via the differential of $\phi$.
(3.49) $T_{m}\mathcal{X}=\ker D\pi[m]\oplus(\ker D\pi[m])^{\perp}=\mathcal{V}_{m}\oplus\mathcal{H}_{m},$
@@
Let $(\mathcal{X},g)$ and $(\mathcal{Y},h)$ be two Riemannian manifolds and let $T_{m}\mathcal{X}=\mathcal{V}_{m}\oplus\mathcal{H}_{m}$ be the direct sum decomposition of the tangent space at $m\in\mathcal{X}$. A mapping $\pi:\mathcal{X}\to\mathcal{Y}$ is a Riemannian submersion if and only if the differential $D\pi[m]:T_{m}\mathcal{X}\to T_{\pi(m)}\mathcal{Y}$ satisfies the following for all $m\in\mathcal{X}$:
@@
A core fact underlying [4, Section 4] and Theorem 2.3 is that understanding the horizontal space of a submersion allows us to identify geodesics that are preserved under the submersion, as described by the following theorem [12, 4].
@@
Let $(\mathcal{X},g)$ and $(\mathcal{Y},h)$ be Riemannian manifolds and $\phi:\mathcal{X}\to\mathcal{Y}$ be a Riemannian submersion. Let $x:[0,1]\to\mathcal{X}$ be a geodesic between $A,B\in\mathcal{X}$ in the geometry $(\mathcal{X},g)$ so that $x^{\prime}(0)\in\mathcal{H}_{A}$ i.e. is horizontal. Then,
@@
    $\phi\circ x$ is a geodesic in $(\mathcal{Y},h)$ between $\phi(A)$ and $\phi(B)$ and is the same length as $x$.
Thus, if we are able to classify horizontal geodesics in $(\mathcal{M},\iota)$, Theorem 3.10 will give explicit formulas for geodesics in $(\mathfrak{M}_{d},g^{N})$ through the submersion $\phi$. The simplest class of geodesics in $(\mathcal{M},\iota)$ are straight lines. Thus, we begin with a proof of Lemma 2.2 that describes a set of straight lines in $\mathcal{M}$.
@@
First, we wish to find the equivalent condition to $\mathbf{W}(t)=(1-t)\mathbf{A}+t\mathbf{B}$ being balanced for all $t\in(0,1)$ when $\mathbf{A},\mathbf{B}\in\mathcal{M}$. Fix $p\in\{1,2,\ldots,N-1\}$. We would like to find an equivalent condition to
@@
where $W_{p}(t)=(1-t)W_{p}(0)+tW_{p}(1)$ for all $p$. Expanding the algebra out and using that $W_{p}(0)=A_{p}$ and $W_{p}(1)=B_{p}$ are both in the balanced manifold, we find that $\mathbf{W}(t)$ is balanced for all $t\in(0,1)$ if and only if
@@
We can rewrite in terms of $M=\tilde{\Lambda}^{-1}R_{p}\Lambda$ to get the final equivalent condition:
(3.54) $R_{p+1}-R_{p-1}$ $=-M(R_{p+1}-R_{p-1})^{T}M^{-T}.$
(3.55) $(R_{p+1}-R_{p-1})M^{T}$ $=-M(R_{p+1}-R_{p-1})^{T}.$
The structure demonstrates that $\mathbf{W}(t)$ is balanced if and only if the difference
(3.56) $(R_{p+1}-R_{p-1})M^{T}$
is a skew symmetric matrix.
It remains to show that if $R_{p}=R_{p-1}$ for all $1\leq p\leq N$, $\mathbf{W}(t)\in\mathcal{M}$ for all $t\in[0,1]$. In this case, the above skew symmetry condition is immediate as all differences $R_{p+1}-R_{p-1}$ are 0, implying that $\mathbf{W}(t)$ is balanced for all $t\in[0,1]$. Thus, we just need to show that the coordinates of $\mathbf{W}(t)$ stay full rank at all intermediate $t$.
Fix $1\leq p\leq N$ and consider the $p$-th coordinate of $\mathbf{W}(t)$. Since $R_{p}$ is the same for all $p$, we drop the index for clarity.
(3.57) $\mathbf{W}_{p}(t)$ $=(1-t)A_{p}+tB_{p}$
(3.58) $=(1-t)Q_{p}\Lambda Q_{p-1}^{T}+t\tilde{Q}_{p}\tilde{\Lambda}\tilde{Q}_{p-1}^{T}$
(3.59) $=(1-t)\tilde{Q}_{p}R\Lambda R^{T}\tilde{Q}_{p-1}^{T}+t\tilde{Q}_{p}\tilde{\Lambda}\tilde{Q}_{p-1}^{T}$
(3.60) $=\tilde{Q}_{p}\left((1-t)R\Lambda R^{T}+t\tilde{\Lambda}\right)\tilde{Q}_{p-1}^{T}.$
Because $\mathbb{P}_{d}$ is convex, the interior convex combination is also in $\mathbb{P}_{d}$ as $R\Lambda R^{T}\in\mathbb{P}_{d}$. Since $\tilde{Q}_{p},\tilde{Q}_{p-1}\in O_{d}$ and are invertible, we conclude that $\mathbf{W}_{p}(t)$ is full rank for all $t\in[0,1]$ and $\mathbf{W}(t)\in\mathcal{M}$ follows. ∎
We now state a basis for the horizontal space in $\mathcal{M}$ computed in [10, Section 6] and [13, Theorem 34]. Let
@@
so that $\mathbf{M}=\xi(\Lambda,Q_{N},Q_{N-1},\ldots,Q_{1})$ and $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{d})$. The horizontal space at $\mathbf{M}$ is $\mathcal{H}_{\mathbf{M}}=\ker(\phi^{*}[\mathbf{M}])^{\perp}$. Let $q_{p,i}$ denote the $i$-th column of $Q_{p}$.
Lemma 3.11 (Basis for Horizontal Space in $\mathcal{M}$).
Let $Z$ be an arbitrary matrix in $\mathbb{M}_{d}^{N}$. Then, $Z\in\mathcal{H}_{\mathbf{M}}$ if and only if there exists a set of coefficients $\alpha\in\mathbb{M}_{d}$ such that
@@
for all $1\leq p\leq N$. Equivalently, $\mathcal{H}_{\mathbf{M}}$ is spanned by
(3.63) $\left\{\left(q_{N,k}q_{N-1,l}^{T},q_{N-1,k}q_{N-2,l}^{T},\ldots q_{1,k}q_{0,l}^{T}\right)|1\leq k,l\leq d\right\}.$
We can now use $\phi$ and Theorem 3.10 to construct new explicit formulas for geodesics on $\mathfrak{M}_{d}$.
@@
where $\xi$ is the map given in Equation 1.6 since $\mathbf{A}\in\mathcal{M}$. Clearly, $\phi(\mathbf{A})=A$. Furthermore, by Lemma 3.11, $\mathbf{Z}\in\mathbb{M}_{d}^{N}$ is in $\mathcal{H}_{\mathbf{A}}$ if and only if there exists $\alpha\in\mathbb{M}_{d}$ such that
@@
It is immediate that $\mathbf{A}$ is in its own horizontal space $\mathcal{H}_{\mathbf{A}}$ using the coefficient matrix $\alpha=\Sigma^{\frac{1}{N}}$.
@@
where $Q=U^{T}\tilde{U}=V^{T}\tilde{V}$ by assumption. Clearly, $\phi(\mathbf{B})=B$. Furthermore, $\mathbf{A}$ and $\mathbf{B}$ satisfy the sufficient condition of Lemma 2.2 since $Q^{T}=\tilde{U}^{T}U=\tilde{V}^{T}V$ and $R_{p}$ is constant for all $p$, so we conclude that $x(t)\in\mathcal{M}$ for all $t\in[0,1]$. Thus, the line $(1-t)\mathbf{A}+t\mathbf{B}$ is indeed the geodesic between $\mathbf{A}$ and $\mathbf{B}$ in $(\mathcal{M},\iota)$ because $\iota$ is flat.
Now we will show $x^{\prime}(0)$ is horizontal for the chosen $\mathbf{A}$ and $\mathbf{B}$. We have already shown that $\mathbf{A}$ is in $\mathcal{H}_{\mathbf{A}}$. By Equation 3.65, it is enough to show that $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$ as well. Again, Lemma 3.11 provides a sufficient system of equations to solve with $\mathbf{Z}=\mathbf{B}$:
@@
(3.73) $\alpha V^{T}$ $=Q\tilde{\Sigma}^{\frac{1}{N}}\tilde{V}^{T}.$
The solution given by the middle equations of $\alpha=Q\tilde{\Sigma}^{\frac{1}{N}}Q^{T}$ can be checked as consistent with the others by definition of $Q$. Thus, $\mathbf{B}\in\mathcal{H}_{\mathbf{A}}$.
Since we have shown that both $\mathbf{B}$ and $\mathbf{A}$ are in $\mathcal{H}_{\mathbf{A}}$ and that the geodesic linking them is a straight line, it follows from Equation 3.65 that $x^{\prime}(0)$ is horizontal. Furthermore, note that because all coordinates in $x(t)\in\mathcal{M}$, $\phi\circ x\in\mathfrak{M}_{d}$ for all $t\in[0,1]$ as $\phi\circ x$ becomes a product of invertible matrices. Applying Theorem 3.10, we conclude that $\phi\circ x$ is a geodesic in $(\mathfrak{M}_{d},g^{N})$ of the same length as $x$ and the explicit formula comes directly from applying $\phi$ to $x$ between our parameterized $\mathbf{A}$ and $\mathbf{B}$. ∎
@@ 4. Open Problems
The main follow up work is (1) solving the general case geodesic equations for arbitrary endpoints and (2) generalizing geodesic results to low rank (rank <d<d) matrices, not just $\mathcal{M}$, as the Riemannian metric and submersion structure exists in all ranks [3, 10, 13]. Once geodesics and distances are understood in the geometries underlying the DLN, one can consider reconnecting them to the vector field given by the gradient flow to make conclusions about training dynamics, similar to [9, Section 3.3].